if mary paints a room in 3hours and john paints it in 4hours. how many hours for mary and john to paint the same room?
Found 2 solutions by solver91311, JimboP1977:
Answer by solver91311(24713)


If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period.

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.



John


Answer by JimboP1977(311)

1/3+1/4 = 1 hour
7/12 room done in 1 hour
so whole of the room done in 1/(7/12) = 12/7 hours = 1 hour 42 mins and 52 seconds